HUBBLE SPACE TELESCOPE DAILY REPORT #3081 PERIOD COVERED: 0000Z (UTC) 03/26/02 - 0000Z (UTC) 03/27/02 Daily Status Report as of 086/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Thirteen Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3 was used 1}.to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. No problems were encountered. 1.2 Completed FGS/1 9234 (Calibrating the Mass-Luminosity Relation at the End of the Main Sequence) Fine Guidance Sensor #1 was used to calibrate the mass-luminosity relation for stars less massive than 0.2 Msun, with special emphasis on objects near the stellar/brown dwarf border. The proposal completed with no reported errors. 1.3 Completed WF/PC-2 8950 (SM3B Cool Down, Contamination Monitor, and Focus Check) The WF/PC-2 was used to monitor the planned cooldown to -88C, followed by UV monitors, focus checks, as well as decons and associated monitors. No problems were noted. 1.4 Completed WF/PC-2 9254 (Cycle 10 CTE Monitor) The WF/PC-2 was used to monitor CTE changes during cycle 10. The proposal completed with no anomalous activity. 1.5 Completed Two Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 1.6 Completed Two Sets of STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 1.7 Completed WF/PC-2 9043 (Cepheid Distances to Early-type Galaxies) The WF/PC-2 was used to continue observations in the HST Key Project on the Extragalactic Distance Scale and the HST project on the "Calibration of Nearby Type Ia Supernovae'' that have greatly improved our knowledge of the Hubble Constant by providing a solid zero point for the Tully- Fisher {TF} relation and Type Ia Supernovae {SNIa}. However, severe inconsistencies remain for distance estimators to early-type galaxies such as surface brightness fluctuations {SBF}, the planetary nebula luminosity function {PNLF}, the fundamental plane {FP}, and the globular cluster luminosity function {GCLF}. As a result, the distance to the Virgo cluster core remains uncertain by as much as 20 determination is directly affected by a lingering 0.1 mag {5 uncertainty in the photometric calibration of the WFPC2. Resolving these issues is essential not only to firm up the extragalactic distance scale, but also to understand the mass and velocity structure of the local universe. SBF in particular is emerging as the method of choice for mapping local velocity fields to 10, 000 kms because it offers an order of magnitude less Malmquist bias than TF, and SNIa are too rare to study large scale flows effectively. This project will tighten the photometric calibration of the WFPC2, and provide a solid Cepheid calibration for SBF and PNLF. The observations completed nominally. 1.8 Completed Two Sets of ACS/WFC/HRC 9013 (Coarse Corrector Alignment) The Advanced Camera for Surveys (WFC and HRC) was used in an attempt to approximately optimize the ACS image quality over the fields of both the WFC and HRC channels by adjusting the IM1 and M1 corrector mechanisms in both tip/tilt {cylinder rotation} and focus. These observations of a moderate density star field constitute the "first light" external images for the HST/ACS. No anomalous activity occurred. 1.9 Completed ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 7 Scheduled Re-acquisitions: 7 Successful: 7 2.2 FHST Updates: Scheduled: 15 Successful: 15 2.3 Operations Notes: The NICMOS Cryo Cooler cooldown continues nominally. Using ROP SR-1A, the SSR-1 EDAC error counter was cleared twice. The following Operations Notes were executed during this operational day: Turn On ACS HRC/SBC Corrector Status limit Adjust ACS Error Count limit Turn Off ACS HRC/SBC Corrector Status limit An SSR Command error was cleared at 085/1248Z per ROP SR-9A. Per an operations request, completed ACS Coarse Corrector Alignment, Part 2 at 085/1738Z. The NSSC-1 status buffer was dumped at 086/0007Z, using ROP NS-3. The Coarse Sun Sensor 3 Temperature (GCSS3T) broke the yellow limit at 086/043437Z and returned to normal bounds at 086/043443Z. During this time the HST was in an SAA interval. The temperature reached -45.0897 degC: the yellow limit is set at -45.0000 degC. HSTAR 8581 was written. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW